OpenStudy (anonymous):

Can someone help me with linear inequalities please! -2/3y>4

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

To get y, multiply by the reciprocal.

OpenStudy (anonymous):

(-2/3)(-3/2)y= 4(-3/2)

OpenStudy (anonymous):

When you multiply or divide by a negative number in an inequality, you must flip the sign. y<-6

OpenStudy (anonymous):

Ok, thank you!! Would it be a problem if you can help me with 2 more problems please.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Ok, the first is \[\frac{ 1 }{ 5 }y-\frac{ 1 }{ 3 } \le \frac{ 2 }{ 3 }\] and the second one is \[\frac{ 2 }{ 3 }x-5 \le 7\]

OpenStudy (anonymous):

1st one You want y by itself so move the -1/3 to the other side.

OpenStudy (anonymous):

Do u know how to do that?

OpenStudy (anonymous):

After u do that you should have 1/5y less than or equal to 1

OpenStudy (anonymous):

Yea, thats what I got. Is that my final answer, because on te back of my book it said the answer would be \[y \le -5\]

OpenStudy (anonymous):

Almost. It's 5 where'd u get the neg

OpenStudy (anonymous):

oh it's because 1/5y is actual -1/5y, Sorry. lol

OpenStudy (anonymous):

Oh. Ok

OpenStudy (anonymous):

Do u know what to do on the next one

OpenStudy (anonymous):

Would be the same one as the first one?

OpenStudy (anonymous):

Sort of give it a shot

OpenStudy (anonymous):

The equation is this right 2/3x−5≤7, so I would add 5 to both sides, and that would get \[\frac{ 2 }{ 3 }x \le 7\], right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

not 7 but 12.

OpenStudy (anonymous):

12

OpenStudy (anonymous):

Then I would divide both side by 2/3, correct?

OpenStudy (anonymous):

Or multiply by the reciprocal 3/2

OpenStudy (anonymous):

ok, thank you! You were a big help =)

OpenStudy (anonymous):

You're welcome